# Language-guided Open-world Video Anomaly Detection

### v1

@@ Abstract
Video anomaly detection models aim to detect anomalies that deviate from what is expected. In open-world scenarios, the expected events may change as requirements change. For example, not wearing a mask is considered abnormal during a flu outbreak but normal otherwise. However, existing methods assume that the definition of anomalies is invariable, and thus are not applicable to the open world. To address this, we propose a novel open-world VAD paradigm with variable definitions, allowing guided detection through user-provided natural language at inference time. This paradigm necessitates establishing a robust mapping from video and textual definition to anomaly score. Therefore, we propose LaGoVAD (Language-guided Open-world VAD), a model that dynamically adapts anomaly definitions through two regularization strategies: diversifying the relative durations of anomalies via dynamic video synthesis, and enhancing feature robustness through contrastive learning with negative mining. Training such adaptable models requires diverse anomaly definitions, but existing datasets typically provide given labels without semantic descriptions. To bridge this gap, we collect PreVAD (Pre-training Video Anomaly Dataset), the largest and most diverse video anomaly dataset to date, featuring 35,279 annotated videos with multi-level category labels and descriptions that explicitly define anomalies. Zero-shot experiments on seven datasets demonstrate SOTA performance. Data and code will be released.
@@ 1 Introduction
Video Anomaly Detection (VAD) aims to identify frames in videos that deviate from expected patterns [5, 33], which is applicable in fields such as intelligent surveillance and monitoring [20]. In recent years, many VAD methods have achieved commendable performance employing weak supervision [21, 14, 37, 6, 25, 35] or semi-supervision [24, 31] in the closed-set setting. However, there is a consensus [34, 20, 46, 45] that the algorithm should detect anomalies beyond the training data in open-world scenarios.
Figure 1: Comparison of different VAD paradigms. Closed-set methods (b) can only detect anomalies in the training scope, while open-set methods (c) can detect novel anomalies. Our open-world approach (d) can deal with changes to the pattern’s label in open-world scenarios, with an example in (e).
@@
As shown in Fig. 1a, the training data for VAD models encompass patterns labeled as normal or abnormal, where normal patterns include activities such as running and abnormal patterns comprise events like explosions. Conventional closed-set methods (Fig. 1b) [21, 35] aim to detect patterns identical to those encountered during training when applied to test sets, thereby restricting their application in open-world scenarios. In contrast, open-set approaches (Fig. 1c) [46] (including open-vocabulary [34] and domain generalization [2, 13, 30] methods) are able to detect novel patterns absent from the training data without tuning. However, these methods neglect the critical issue of potential label alteration during testing (Fig. 1d), i.e., patterns originally labeled as normal may be redefined as abnormal (and vice versa). A representative example from Fig. 1e demonstrates this phenomenon: while pedestrian on road is regarded as a normal behavior in conventional crime anomaly datasets [25], this same pattern would typically be classified as abnormal in freeway surveillance scenarios. The cause of such label alteration lies in the user’s different definition of what constitutes anomalies, driven by environments or temporal policies. Formally, this is a concept drift issue, as defined in [18], which refers to the divergence between the conditional probability distributions of training and testing phases, i.e., $P_{\text{train}}(y|v)\neq P_{\text{test}}(y|v)$, where $v$ are videos and $y$ are anomaly labels. While some attempts have begun to address this, critical limitations remain. Scene-dependent methods [4, 8, 2] associate the anomaly definition with scenes, neglecting user-specific requirements (e.g., hospital administrators may require detecting the anomaly of not wearing masks during influenza outbreaks but not at other times). Meanwhile, a dataset-dependent method [9] explores anomaly conflicts across datasets, but remains constrained by predefined categories of training datasets, lacking generalizability to open-world dynamics. Besides the limitation of task paradigms, existing methods are evaluated only on limited scenes with small-scale data, lacking extensive zero-shot cross-domain comparisons to verify the open-world capability.
@@
To address the aforementioned concept drift challenge, we propose a novel open-world paradigm. First, we explicitly model the anomaly definition as a stochastic variable instead of fixing it as one or a few realizations. Then, we condition predictions $y$ on both the video $v$ and the anomaly definition $z$, i.e., learning a mapping $\Phi:(v,z)\rightarrow y$. Since we take the changing definition of anomalies into account, we effectively avoid concept drift (as detailed in Sec. 3). Finally, to enable natural interaction, we employ textual anomaly definition, allowing users to dynamically define anomalies via language.
@@
However, learning $\Phi$ requires modeling a more complex multimodal space, resulting in decaying sample density that leads to overfitting. To address this, we mitigate it from both model and dataset perspectives. We propose a Language-Guided Open-world Video Anomaly Detector (LaGoVAD), which employs two regularization strategies to reduce overfitting: 1) We align vision and language through contrastive learning with negative sample mining, which helps the model learn more robust features. 2) We incorporate a novel dynamic video synthesis module that generates long videos and pseudo-labels on the fly, which functions as a form of data augmentation to diversify the relative duration of abnormal events.
We construct a large-scale diversified Pre-training Video Anomaly Dataset (PreVAD), which is collected through a scalable data curation pipeline utilizing foundation models to automate data cleaning and annotation, significantly reducing manual labeling costs while ensuring high quality. PreVAD comprises 35,279 videos annotated with multi-level categories and anomaly descriptions. To our knowledge, PreVAD surpasses existing datasets in diversity and scale. To evaluate the open-world capability, we set up a new benchmark that performs zero-shot anomaly detection on seven VAD datasets (UCF-Crime [25], XD-Violence [32], MSAD [45], UBNormal [1], DoTA [38], TAD [17], LAD [27]). Our contributions are summarized as follows:
@@
    We reformulate open-world VAD that pioneers the formulation of the concept drift in VAD and proposes a language-guided paradigm to avoid it.
@@
    We conduct zero-shot evaluation across seven datasets to validate the generalization of our method, and LaGoVAD achieves state-of-the-art performance (Fig. 2).
Figure 2: Comparison of zero-shot performance between LaGoVAD and other methods across seven datasets [25, 32, 1, 38, 27, 17, 45]. Except for XD-Violence which uses the AP as metric, all other datasets adopt the AUC metric for evaluation. LaGoVAD significantly outperforms existing state-of-the-art methods on all datasets.
@@ 2 Related Work
Table 1: Comparisons between PreVAD and existing datasets. Our dataset 1) has the largest scale and broadest domain coverage, 2) is annotated with abnormal video descriptions, 3) does not integrate data from existing VAD datasets.
@@
Domain # categories Text Annotation Source
ShanghaiTech [42] 437 (107) campus 13 - recording
UCF-Crime [25] 1900 (950) crime 14 - web
XD-Violence [32] 4754 (2405) crime 7 - web,movie
UBI-Fights [10] 1000 (216) violence 2 - web
LAD [27] 2000 (762) crime, traffic, animal, mishap 14 - web
TAD [17] 500 (250) crime 8 - web
UBNormal [1] 543 (278) pedestrian 28 - synthesis
DoTA [38] 5677 (5677) traffic 9 - web
NWPU-Campus [4] 547 (124) campus 28 - recording
MSAD [45] 720 (240) crime, traffic, mishap 55 - web
UCCD [43] 1012 (382) crime - dense UCF
UCA [39] 1854 (944) crime - dense UCF
VAD-Instruct50k [41] 5547 (2715) crime - instruction UCF+XD
HAWK [26] 7852 (6677) crime, traffic - instruction 7 VAD datasets
CUVA [12] 1000 (1000) crime, traffic, pedestrian, animal 42 instruction web
PreVAD 35279 (11979) crime, traffic, animal, mishap, production 35 anomaly description web
2.1 Video Anomaly Datasets
We summarize the characteristics of existing video anomaly datasets in Tab. 1.
Scale. The largest standalone dataset [32] contains only 5K videos, with ensemble datasets reaching 7.8K [26]. The data scarcity limits the performance of VAD.
Domain & Category. Many datasets focus only on a single scene, such as traffic or campus. The few datasets that cover multiple scenes overlook domains like mishaps, animal-related violence, and production accidents.
Text Annotation. Existing VAD datasets are labeled with anomaly categories, which introduces semantic ambiguity. Although [43, 39, 41, 12, 26] provide different types of text annotation, they focus on understanding or captioning tasks and cannot provide a fine-grained overall description of the anomaly in a video.
Source. Current datasets are mainly from public web videos, while others rely on synthetic generation [1, 19, 34] or movie clips [34]. However, synthetic datasets suffer from misalignment with the real world, and movie data raises concerns about potential copyright infringement.
In this paper, we propose a scalable data curation pipeline to collect a novel dataset, which has large-scale diversified videos with multi-level taxonomy and anomaly descriptions.
@@ 2.2 Open-world Video Anomaly Detection Methods
Intuitively, open-world VAD models should detect novel anomalies beyond the training set, as discussed in [46, 34, 26, 13]. From a task paradigm perspective, early attempts adopt open-set and domain generalization strategies [1, 46, 13]. Then, [34] extends this paradigm with open-vocabulary VAD, enabling both detection and classification of unseen anomalies. However, these approaches implicitly assume a fixed anomaly definition and restrict model exposure to partial categories during training, unable to deal with the concept drift issue. Recent studies explore the dynamic anomaly definition: [4, 8, 2] posit that anomaly is scene-dependent (e.g., identical behaviors classified differently across scenes), training models to infer scene-anomaly correlations from data, and [9] trains dataset-specific classifiers. Despite these efforts, they lack the ability of user-customizable anomaly definition, limiting their applicability in open-world scenarios. Additionally, [26] explores open-world video understanding for video-QA tasks but lacks detection capabilities.
From a model design perspective, current advancements primarily adopt two pathways: 1) data-driven approaches [1, 46, 13, 34, 19] enhance generalization by utilizing more data, while 2) cross-modal alignment approaches [6, 35, 34, 37] aim to construct more robust feature spaces by aligning vision and language. However, they neglect the problem of duration distribution shifts when leveraging more data and only align videos to class-level text embeddings without further fine-grained aligning.
@@ 3 Paradigm: Language-guided Open-world Video Anomaly Detection
Figure 3: Architecture of our proposed LaGoVAD. We implement the language-guided VAD paradigm by adding an anomaly definition branch ($z\rightarrow\mathcal{G}\rightarrow\mathcal{U}$). The model is trained with two novel regularization strategies: dynamic video synthesis $\mathcal{L}_{\text{dvs}}$ (Sec. 4.1) and contrastive learning loss with negative mining $\mathcal{L}_{\text{neg}}$ (Sec. 4.2).
We define open-world video anomaly detection as the task of identifying video frames containing abnormal patterns, where the definition of abnormality may change during testing. Abnormal patterns manifest as events, behaviors, or actions (e.g., running). In practice, the definition of anomalies may change as users’ requirements change, influenced by cultural differences, policy updates, and specific environments. The user may expand the definition to detect new anomalies or narrow the definition to remove those of no interest, which causes the abnormality label of a particular pattern to change. For instance, while running is generally normal behavior, it becomes abnormal in libraries or offices. Formally, this is a concept drift issue [18]:
@@
where $V$ denotes the video and $Y$ denotes the label. This conditional probability distribution can be expanded as:
@@
where $Z$ denotes the anomaly definition. We hypothesize that $Y$ is solely determined by $Z$ and $V$, i.e., the result only depends on the video to be detected and the anomaly definition. Therefore, the concept drift happens due to the change of $P(Z|V)$.
@@
where $\theta$ denotes the parameters of the model $\Phi$, and $\mathcal{L}$ denotes the loss function. It is worth emphasizing that some methods that can detect unknown anomalies also belong to this paradigm, including open-set [46, 1], domain generalization [30] and open-vocabulary [34] methods, because they assume a fixed category set under a specific definition and only a subset are available in training. Under their assumption, an abnormal pattern would never change to normal, and thus they are unable to deal with the concept drift in open world.
@@
During training, the model $\Phi$ learns an optimal group of parameters $\theta$ that detect anomalies in video $v$ under the guidance of definition $z$. We implement $z$ in the form of natural language, but theoretically, it could be image, video, audio, or a learned embedding. Our paradigm is especially applicable when the user needs to temporarily specify an abnormal behavior, e.g., detecting not wearing a mask during a flu outbreak.

### v2

@@ Abstract
Video anomaly detection (VAD) aims to detect anomalies that deviate from what is expected. In open-world scenarios, the expected events may change as requirements change. For example, not wearing a mask may be considered abnormal during a flu outbreak but normal otherwise. However, existing methods assume that the definition of anomalies is invariable, and thus are not applicable to the open world. To address this, we propose a novel open-world VAD paradigm with variable definitions, allowing guided detection through user-provided natural language at inference time. This paradigm necessitates establishing a robust mapping from video and textual definition to anomaly scores. Therefore, we propose LaGoVAD (Language-guided Open-world Video Anomaly Detector), a model that dynamically adapts anomaly definitions under weak supervision with two regularization strategies: diversifying the relative durations of anomalies via dynamic video synthesis, and enhancing feature robustness through contrastive learning with negative mining. Training such adaptable models requires diverse anomaly definitions, but existing datasets typically provide labels without semantic descriptions. To bridge this gap, we collect PreVAD (Pre-training Video Anomaly Dataset), the largest and most diverse video anomaly dataset to date, featuring 35,279 annotated videos with multi-level category labels and descriptions that explicitly define anomalies. Zero-shot experiments on seven datasets demonstrate LaGoVAD’s SOTA performance. Our dataset and code will be released at https://github.com/Kamino666/LaGoVAD-PreVAD.
@@ 1 Introduction
Video Anomaly Detection (VAD) aims to identify frames in videos that deviate from expected patterns (Chandrakala et al., 2023; Wu et al., 2024a), which is applicable in fields such as intelligent surveillance (Pang et al., 2020). In recent years, many VAD methods have achieved commendable performance employing weak supervision (Pu et al., 2024; Joo et al., 2023; Yang et al., 2024b; Chen et al., 2023; Sultani et al., 2018; Wu et al., 2024c) in the closed-set setting. However, there is a consensus (Wu et al., 2024b; Pang et al., 2020; Zhu et al., 2022; 2024b) that the field is moving towards enabling models to detect anomalies beyond the training data in open-world scenarios.
@@
As shown in Fig. 1a, the training data of VAD models encompass patterns labeled as normal or abnormal, where normal patterns include activities such as running or driving and abnormal patterns include events like explosions. Conventional closed-set methods (Fig. 1b) (Pu et al., 2024; Wu et al., 2024c) aim to detect patterns identical to those encountered during training when applied to test sets, thereby restricting their application in open-world scenarios. In contrast, open-set approaches (Fig. 1c) (Zhu et al., 2022) (including open-vocabulary (Wu et al., 2024b) and domain generalization (Aich et al., 2023; Jain et al., 2024; Wang et al., 2024c) methods) are able to detect novel patterns absent from the training data without tuning. However, these methods neglect the critical issue of potential label change during testing (Fig.1d), i.e., patterns originally labeled as normal may be redefined as abnormal (and vice versa). A representative example from Fig. 1e demonstrates this phenomenon: while pedestrian on road is regarded as a normal behavior in conventional crime anomaly datasets, this same pattern would typically be classified as abnormal in freeway surveillance scenarios. The cause of such label change lies in the user’s different definition of what constitutes anomalies, driven by environments or temporal policies. Formally, this is a concept drift issue, as defined in (Moreno-Torres et al., 2012), which refers to the divergence between the conditional probability distributions of training and testing phases, i.e., $P_{\text{train}}(Y|V)\neq P_{\text{test}}(Y|V)$, where $V$ are videos and $Y$ are anomaly labels.
Figure 1: Comparison of different VAD paradigms. Closed-set methods (b) can only detect anomalies in the training scope, while open-set methods (c) can detect novel anomalies. Our open-world approach (d) can deal with label change in open-world scenarios, with an example in (e).
While some attempts have begun to address this, critical limitations remain. Scene-dependent methods (Cao et al., 2023; Cho et al., 2023; Aich et al., 2023) associate the anomaly definition with scenes, neglecting user-specific requirements (e.g., hospital administrators may require detecting the anomaly of not wearing masks during influenza outbreaks but not at other times). Meanwhile, a dataset-dependent method (Cho et al., 2024) explores anomaly conflicts across datasets, but remains constrained by predefined categories of training datasets. Besides the limitation of task settings, existing methods are evaluated on limited scenes with small-scale data, lacking extensive zero-shot cross-domain comparisons to verify the open-world capability.
@@
To address the concept drift challenge, we propose a novel open-world paradigm. First, we explicitly model the anomaly definition as a stochastic variable instead of fixing it as one or a few realizations. Then, we condition predictions $Y$ on both the video $v$ and the anomaly definition $Z$, i.e., learning a mapping $\Phi:(V,Z)\rightarrow Y$. Since we take the changing definition into account, we effectively avoid concept drift (as detailed in Sec. 3). Finally, to enable natural interaction, we employ textual anomaly definition, allowing users to dynamically define anomalies via language.
@@
However, learning $\Phi$ requires modeling a more complex multimodal space, resulting in decaying sample density that leads to overfitting. To address this, we mitigate it from both model and dataset perspectives. As for the model, we propose a Language-Guided Open-world Video Anomaly Detector (LaGoVAD), which employs two regularization strategies to reduce overfitting: 1) Aligning vision and language through contrastive learning with negative sample mining, which learns more robust features. 2) Incorporating a dynamic video synthesis module that generates long videos and pseudo-labels on the fly, which diversifys the relative duration of abnormal events. As for the dataset, we construct a large-scale diversified Pre-training Video Anomaly Dataset (PreVAD), collected through a scalable data curation pipeline utilizing foundation models to automate data cleaning and annotation, significantly reducing manual labeling costs while ensuring high quality. PreVAD comprises 35,279 videos annotated with multi-level categories and anomaly descriptions, supporting weakly-supervised training. To our knowledge, PreVAD surpasses existing datasets in diversity and scale.
The model is evaluated under two zero-shot evaluation protocols: one comprehensively assesses open-world capability by evaluating cross-domain performance across seven diverse datasets, addressing key open-world challenges such as detecting unseen categories and handling concept drift, while the other specifically measures concept drift by averaging performance on a dataset under different anomaly definitions. Our contributions are summarized as follows:
@@
    We reformulate open-world VAD that pioneers the formulation of the concept drift in VAD and proposes a joint modeling paradigm to avoid it.
@@
    We conduct zero-shot cross-dataset evaluation and concept drift evaluation to validate the generalization, where LaGoVAD achieves state-of-the-art performance.
@@ 2 Related Work
2.1 Video Anomaly Datasets
We summarize the characteristics of existing video anomaly datasets in Tab. 1. Scale: The largest standalone dataset (Wu et al., 2020) contains only 5K videos, with ensemble datasets reaching 7.8K (Tang et al., 2024). The data scarcity limits the performance of VAD. Domain & Category: Many datasets focus only on a single scene, such as traffic or campus. The few datasets that cover multiple scenes overlook domains like mishaps, animal-related violence, and production accidents. Text Annotation: Existing VAD datasets are labeled with anomaly categories, which introduces semantic ambiguity. Although some datasets provide different types of text annotation, they focus on understanding or captioning tasks and cannot provide a fine-grained overall description of the anomaly in a video. Source: Current datasets are mainly from public web videos, while others rely on synthetic generation (Acsintoae et al., 2022; Narwade et al., 2024) or movie clips (Wu et al., 2020). However, synthetic datasets suffer from misalignment with the real world, and movie data raises concerns about potential copyright infringement. In this paper, we propose a scalable data curation pipeline to collect a novel dataset, which has large-scale diversified videos with multi-level taxonomy and anomaly descriptions.
Table 1: Comparisons between PreVAD and existing datasets. Our dataset 1) has the largest scale and broadest domain coverage, 2) is annotated with abnormal video descriptions, 3) enables zero-shot evaluation without relying on existing VAD datasets.
@@ 2.1 Video Anomaly Datasets
Domain # categories Text Anno. Source
ShanghaiTech (Zhong et al., 2019) 437 (107) campus 13 - recording
UCF-Crime (Sultani et al., 2018) 1900 (950) crime 14 - web
XD-Violence (Wu et al., 2020) 4754 (2405) crime 7 - web,movie
LAD (Wan et al., 2021) 2000 (762) crime, traffic, animal, mishap 14 - web
TAD (Lv et al., 2021) 500 (250) crime 8 - web
UBNormal (Acsintoae et al., 2022) 543 (278) pedestrian 28 - synthesis
DoTA (Yao et al., 2022) 5677 (5677) traffic 9 - web
MSAD (Zhu et al., 2024b) 720 (240) crime, traffic, mishap 55 - web
UCCD (Zhou et al., 2024) 1012 (382) crime - dense UCF
UCA (Yuan et al., 2024) 1854 (944) crime - dense UCF
VAD-Instruct50k (Zhang et al., 2024) 5547 (2715) crime - instruction UCF+XD
HAWK (Tang et al., 2024) 7852 (6677) crime, traffic - instruction 7 VAD datasets
CUVA (Du et al., 2024) 1000 (1000) crime, traffic, pedestrian, animal 42 instruction web
PreVAD 35279 (11979)
crime, traffic, animal,
mishap, production
35
anomaly
description
web
@@ 2.2 Open-world Video Anomaly Detection Methods
Intuitively, open-world VAD models should detect novel anomalies beyond the training set (Zhu et al., 2022; Wu et al., 2024b; Tang et al., 2024; Jain et al., 2024). From a task paradigm perspective, early attempts adopt open-set and domain generalization strategies (Acsintoae et al., 2022; Zhu et al., 2022; Jain et al., 2024). Then, Wu et al. (2024b) extends this paradigm with open-vocabulary VAD, enabling both detection and classification of unseen anomalies. However, these approaches implicitly assume a fixed anomaly definition and restrict model exposure to partial categories during training, unable to deal with the concept drift issue. Recent studies explore the dynamic anomaly definition: Cao et al. (2023); Cho et al. (2023); Aich et al. (2023) posit that anomaly is scene-dependent (e.g., identical behaviors classified differently across scenes), training models to infer scene-anomaly correlations from data, and Cho et al. (2024) trains dataset-specific classifiers. Despite these efforts, they lack the ability of user-customizable anomaly definition, limiting their applicability in open-world scenarios. Additionally, Tang et al. (2024) explores open-world video understanding for video-QA tasks but lacks detection capabilities. And although Yang et al. (2024a) incorporates an induction-deduction strategy to detect anomalies based on rules extracted from few samples, it is limited to one-class and few-shot contexts.
From a model design perspective, current advancements primarily adopt two ways: 1) data-driven approaches (Acsintoae et al., 2022; Zhu et al., 2022; Jain et al., 2024; Wu et al., 2024b) enhance generalization by utilizing more data, while 2) cross-modal alignment approaches (Chen et al., 2023; Wu et al., 2024c; Yang et al., 2024b) aim to construct more robust feature spaces by aligning vision and language. However, they neglect the problem of duration distribution shifts when leveraging more data and only align videos to class-level text embeddings without further fine-grained aligning.
@@ 3 Paradigm: Language-guided Open-world Video Anomaly Detection
We define open-world video anomaly detection as the task of identifying video frames containing abnormal patterns, where the definition of abnormality may change during testing. Abnormal patterns manifest as events, behaviors, or actions (e.g., running). In practice, the definition of anomalies may change as requirements change, influenced by cultural differences, policy updates, and specific environments. The user may expand the definition to detect new anomalies or narrow the definition to remove those of no interest, which causes the label of a particular pattern to change. For instance, while running is generally normal behavior, it becomes abnormal in libraries or offices. Formally, this is a concept drift issue (Moreno-Torres et al., 2012):
@@
where $V$ denotes the video and $Y$ denotes the label, which can be expanded as:
@@
where $Z$ denotes the anomaly definition. We hypothesize that $Y$ is solely determined by both $Z$ and $V$, i.e., the result depends on both the video and the anomaly definition. Therefore, the concept drift happens due to the change of $P(Z|V)$.
@@
where $\theta$ denotes the parameters of the model $\Phi$, and $\mathcal{L}$ denotes the loss function. It is worth emphasizing that some methods that can detect unknown anomalies also belong to this paradigm, including open-set (Zhu et al., 2022; Acsintoae et al., 2022), domain generalization (Wang et al., 2024c) and open-vocabulary (Wu et al., 2024b) methods, because they assume a fixed category set under a specific definition and only a subset are available in training. Under their assumption, an abnormal pattern would never change to normal, and thus they are unable to deal with the concept drift.
@@
During training, the model $\Phi$ learns an optimal group of parameters $\theta$ that detect anomalies in video $v$ under the guidance of definition $z$. We later implement $z$ in the form of natural language, but theoretically, it could be images, videos, audio, or a learned embedding.
Figure 2: Architecture of our proposed LaGoVAD, which implement Eq. 4 by adding an anomaly definition branch ($z\rightarrow\mathcal{G}\rightarrow\mathcal{U}$). The model is trained with two novel regularization strategies: dynamic video synthesis $\mathcal{L}_{\text{dvs}}$ (4.1) and contrastive learning loss with negative mining $\mathcal{L}_{\text{neg}}$ (4.2).